OpenStudy (anonymous):

(a+2/15)-(a+5/5)<(a+10/30)

OpenStudy (anonymous):

\[\frac{a+2}{15}-\frac{a+5}{5} <\frac{a+10}{30}\] multiply both sides by 30 to get \[2(a+2)-6(a+5)<a+10\] and should be good from there

OpenStudy (anonymous):

just don't forget the parentheses and the distributive law. you get \[2a+4-6a+30<a+10\] \[-4a+34<a+10\] \[34-10<a+4a\] \[24<5a\] \[\frac{24}{5}<a\] if my algebra is correct

OpenStudy (anonymous):

ok no my algebra is wrong

OpenStudy (anonymous):

\[2a+4-6a-30 < a+10\] \[-4a-26<a+10\] \[-26-10<a+4a\] \[-36<5a\] \[-\frac{36}{5}<a\]

OpenStudy (anonymous):

forgot to distribute the "-6" correctly